The RW, Inc. is proposing an investment in energy-saving equipment. The new equipment requires a $9 million investment. The attraction of the new equipment is that it would create an operating cost savings by cutting manufacturing costs. Currently, these costs are $8 per unit. However, as the following table shows, there is some uncertainty both about future sales and about the performance of the new machinery: Pessimistic Expected Optimistic Sales (Millions of Cost with new Economic life of new 0.4 0.5 4 10 0.7 3 13 Units) machinery (S/Unit) 7 machine (Years) Management feels that the new equipment would be so customized to RWs needs that there would be no used equipment market for it, and consequently, it would have a zero salvage value. The cost of capital is 12%. Assume working capital would be unchanged and RW pays no taxes. Under each scenario, do the operating cost savings justify the investment in the new equipment? Hint: Calculate the annual operating cost savings net of the EAC of the initial investment for each scenario a. Pessimistic: No: Expected: No: Optimistic: No b. Pessimistic: No; Expected: No; Optimistic: Yes c. Pessimistic: No; Expected: Yes; Optimistic: Yes d. Pessimistic: Yes; Expected: Yes: Optimistic: Yes
Answer #1
Pessimistic Expected Optimistic
Sales in units 400000 500000 700000
Current cost per unit ($) 6 4 3
Cost per unit with new machinery ($) 8 8 8
Costs savings per unit if, new machinery is istalled 2 4 5
Total cost savings per year 800000 2000000 3500000
EAC of new machine:
=-9000000*0.12*1.12^7/(1.12^7-1) = -1972060
=-9000000*0.12*1.12^10/(1.12^10-1) = -1592857
=-9000000*0.12*1.12^13/(1.12^13-1) = -1401095
Equivalent Annual Present Worth -1172060 407143 2098905
Amswer:
Option [c] = No Yes Yes